# On Marginal Stability in Low Temperature Spherical Spin Glasses

### v2

@@ 1 Introduction
${\mathcal{K}}=\left\{(\zeta,L)\in{\mathcal{N}}\times(0,\infty)\leavevmode\nobreak\ :\leavevmode\nobreak\ L>\int_{0}^{1}\zeta(s)\leavevmode\nobreak\ {\rm d}s\right\}.$
@@
${\mathcal{Q}}(\zeta,L)=\frac{1}{2}\left(\xi^{\prime}(0)L+\int_{0}^{1}\xi^{\prime\prime}(q)\hat{\zeta}(q)\leavevmode\nobreak\ {\rm d}q+\int_{0}^{1}\frac{{\rm d}q}{\hat{\zeta}(q)}\right)=\frac{1}{2}\left(\xi^{\prime}(1)L+\int_{0}^{1}\xi^{\prime\prime}(q)\Big{(}\int_{0}^{q}\zeta(s){\rm d}s\Big{)}\leavevmode\nobreak\ {\rm d}q+\int_{0}^{1}\frac{{\rm d}q}{\hat{\zeta}(q)}\right).$
@@ 1.1 Main Results
For any $\xi$ and $\varepsilon>0$, there exists $\delta>0$ such that the following holds with probability $1-e^{-cN}$ for large $N$. For all $\delta$-approximate ground states ${\boldsymbol{\sigma}}\in A_{\delta}$, the gradient satisfies
@@
This result is stated as a proposition because a more abstract formula for $\nabla H_{N}({\boldsymbol{\sigma}})$, expressed as a derivative of ${\mathrm{GS}}(\xi)$, is essentially in [CS17, Remark 2] (or see [Sub24, Corollary 7]). Although we provide a detailed proof for completeness, the only novelty is an integration by parts using the stationarity conditions for $(\zeta,L)$, which yields the more tractable (1.8). Regarding the bulk spectrum (1.9), known concentration estimates imply that uniformly for all ${\boldsymbol{\sigma}}\in{\mathcal{S}}_{N}$, the spectral measure of $\nabla_{\mathrm{sp}}^{2}H_{N}({\boldsymbol{\sigma}})$ is approximately a semi-circle density with radius $2\sqrt{\xi^{\prime\prime}(1)}$ and shifted by the rescaled radial derivative $\partial_{{\mathrm{rad}}}H_{N}({\boldsymbol{\sigma}})=R({\boldsymbol{\sigma}},\nabla H_{N}({\boldsymbol{\sigma}}))$; see the proof of [Sub21, Lemma 3] (which is rephrased below as Lemma 1.11). Thus (1.9) follows routinely from (1.8).
@@
We say $\xi$ exhibits full RSB endpoint behavior when this condition holds. As explained in Corollary 1.14, this condition is implied by (and is presumably generically equivalent to) the more standard definition of full RSB that $\zeta$ is strictly increasing in a neighborhood of $1$. When (1.10) holds, Proposition 1.2 states that the Hessian at any approximate ground state has many near-zero eigenvalues, i.e is marginally stable. Conversely when $\lambda_{+}<0$, it is natural to hope that $H_{N}$ is locally uniformly concave at near-extrema. However it is not obvious why the Hessian does not have outlier eigenvalues. Our next result proves exactly this when $\xi$ contains only even degree terms.111 Theorem 1.3 and Corollary 1.4 extend with no changes when $\gamma_{1}>0$, but do require $\gamma_{p}=0$ for odd $p\geq 3$. The same holds for Corollaries 1.5 and 1.6. However deducing slow mixing from Corollary 1.6 requires modification if $\gamma_{p}>0$, as does Corollary 1.7. Indeed for the former, one must exclude the topologically trivial phase where low-temperature Langevin dynamics does mix rapidly [HS23c, Theorem 1.8]. As discussed further in Subsection 1.4, the idea is to connect extreme Hessian eigenvalues to the ground state energy of certain vector spin glasses, which can be determined to first-order using precise interpolation bounds.
@@
Additionally, the local uniform concavity in (1.14) easily implies a qualitative description of the deep level sets $A_{\delta}$. As stated in Corollary 1.5 below, all connected components $A_{\delta}$ consist of small separated clusters; this is a more geometric formulation of non-marginal stability. A similar result is deduced in [BSZ20, Proposition 9.1] for a special class of $\xi$, for which a version of (1.14) follows from Kac–Rice asymptotics. (The proof in our setting is strictly easier, since the input result in [BSZ20] gives Hessian control only at critical points.)
@@
We pause here to emphasize two points. First, Corollary 1.5 relies crucially on the absence of Hessian outliers stated in (1.11): even a single near-zero outlier eigenvalue could result in $A_{\delta}$ having large components. Second, the full RSB assumption is essentially necessary: when $\lambda_{+}=0$, Corollary 1.5 evidently does not hold at any ${\boldsymbol{\sigma}}\in A_{\delta}$. Indeed, it is predicted in e.g. [CK94, Section 9] that under full RSB behavior, $\delta=O(1/N)$ may suffice for $A_{\delta}$ to have components of macroscopic diameter. Rigorously, it follows (e.g. by generic perturbations as in [Pan13b, Chapter 3.7]) that when $\xi$ is full RSB near overlap $1$ in the sense that $\zeta$ from (1.4) is strictly increasing on $[q,1]$, then for any $\delta>0$ and sufficiently large $N$, with high probability there exist ${\boldsymbol{\sigma}},{\boldsymbol{\sigma}}^{\prime}\in A_{\delta}$ with $R({\boldsymbol{\sigma}},{\boldsymbol{\sigma}}^{\prime})=q$.
Corollary 1.5 yields further consequences for low temperature Gibbs measures in the absence of marginal stability. Using the separation of distinct components, and that low temperature Gibbs samples are approximate ground states, one obtains a plateau property for the autocorrelation function of stationary Langevin dynamics at low temperature. Given an initialization ${\boldsymbol{x}}_{0}\in{\mathcal{S}}_{N}$, the spherical Langevin dynamics are defined by
@@
where $P_{{\boldsymbol{x}}_{t}}^{\perp}=I_{N}-{\boldsymbol{x}}_{t}^{\otimes 2}/N$ is a rank $1$ projection and $\boldsymbol{B}_{t}$ is standard $N$-dimensional Browian motion. It is well known that these dynamics remain on ${\mathcal{S}}_{N}$ almost surely, with unique stationary distribution $\mu_{\beta}$. We omit the proof of the corollary below, which is exactly the same as [AMS23a, Corollary 2.6]. A byproduct is exponentially slow mixing of Langevin dynamics under the same conditions (since $\mu_{\beta}$ is origin-symmetric when $\xi$ is even).
@@
The same $1/N$ scaling of $\varepsilon_{N}$ was identified for pure spherical spin glasses at very low temperature [Sub17b], and established later within the shattered phase by [AMS23a]. Indeed as we explain in Subsection 2.4, Corollary 1.7 follows by the arguments in [AMS23a, Section 5.1]. We note the decay condition on $\varepsilon_{N}$ is sharp for any $(\xi,\beta)$: [AMS23a, Proposition 2.8] shows if $\varepsilon_{N}\leq o(1/N)$, then the expected total variation distance between $\mu_{\beta,0}$ and $\mu_{\beta,\varepsilon_{N}}$ tends to 00; thus (1.16) does not hold, i.e. there is no chaos. We refer to [AMS23a, Section 2.3] for further discussion, and comparison with the overlap-based notion of disorder chaos appearing in e.g. [Cha09, CS17, CHL18, Eld20].
@@ 1.2 Connections to the Kac–Rice Formula
Whenever the Kac–Rice formula suffices to identify the ground state, the radial derivative (and hence bulk Hessian spectrum) at approximate ground states can be read off from the calculation, and the Hessians will have no outliers in the sense of Theorem 1.3. The reason is that the expected number of local maxima at the ground state energy level without the correct radial derivative (or with outlier eigenvalues) is exponentially small. A similar phenomenon holds for topologically trivial spin glasses: those whose external fields are so strong that the landscape has only two critical points, which are the global extrema [BČNS22, XZ22, HS23c]. Topological trivialization is equivalent to the minimizing $\zeta$ being the constant function 00. In general, $\zeta$ can be much more complicated; see [AZ19, AZ22b] for examples. Aside from $1$-RSB and topologically trivial models, the annealed predictions from the Kac–Rice formula are incorrect, and do not suffice to understand the behavior near the ground state (though see [BJ24] for an application at positive temperature). In these cases, physicists have used the replica method to predict the correct quenched critical point counts; see the recent work [FL18, KK23] as well as [BM80, CS95]. These arguments seem difficult to make rigorous at present. Despite these challenges, it is still natural to believe that exponentially rare behavior does not occur at approximate ground states, because the sets $A_{\delta}$ must be rather small. Theorem 1.3 confirms this intuition for the absence of Hessian outliers, but uses a completely different proof technique.
@@
where $\alpha=\sqrt{\xi^{\prime\prime}+\xi^{\prime}-(\xi^{\prime})^{2}}$ and $\xi,\xi^{\prime},\xi^{\prime\prime}$ are evaluated at $1$. Of course, the exact ground state is a critical point of $H_{N}$. When $\xi$ has full RSB endpoint behavior, Proposition 1.2 implies it is marginally stable. Thus one obtains the following inequality, where the full RSB condition is necessary because ${\mathrm{GS}}(\xi)>E_{\infty}$ for pure models. By contrast ${\mathrm{GS}}(\xi)\geq E_{\infty}^{-}(\xi)$ holds for all $\xi$; see [AB13], or [HS23c, Section 7] for extensions.
@@ 1.3 On Marginal Stability
See also [MW15, FU22]. It is additionally believed that “reasonable” optimization algorithms in high-dimension rapidly reach and then get stuck in the “manifold” of marginal states [CK94, Ken24]; indeed [BAKZ22] recently conjectured that finding stable (i.e. non-marginal) local optima is computationally intractable in many disordered systems. As rigorous evidence for this, [HS24a] analyzed the optimal message-passing algorithms to optimize $H_{N}$ over ${\mathcal{S}}_{N}$ and showed the resulting outputs are marginally stable outside of the topologically trivial phase, even in the more general multi-species setting. Additionally, optimal stable algorithms to optimize mean-field spin glass Hamiltonians are now understood to be closely related to full RSB [Sub21, Mon21, AMS21, HS24b, AMS23b, HS23b, MZ24, JSS24]. Corollary 1.4 gives another rigorous link between full RSB and marginal stability, which however concerns low temperature statics rather than efficient algorithms.
@@ 1.4 Proof Ideas
The proofs of (1.11) and (1.12), which ensure the absence of Hessian outliers, are more delicate and given in Subsections 2.2 and 2.3 respectively. Here we perturb $H_{N}$ by considering augmented Hamiltonians of two or three replicas with constrained overlaps very close to $1$. By construction, the near-extrema for the augmented systems are obtained by starting from ${\boldsymbol{\sigma}}\in A_{\delta}$, and perturbing ${\boldsymbol{\sigma}}$ in opposite directions along a Hessian eigenvector. Thus in each case, precise estimates on the augmented ground state energy translate to bounds on the extreme eigenvalues at any approximate ground state. We find explicit interpolation parameters in the corresponding vector Parisi formulas, and perform a first-order Taylor expansion of the resulting bounds for the augmented ground state energy. This yields eigenvalue estimates which turn out to match the edge of the bulk spectrum from (1.9).
Finally Section 3 gives an alternate proof that Gibbs samples are marginally stable if $\beta\to\infty$ slowly with $N$, if $\xi$ is generic in addition to exhibiting full RSB endpoint behavior. Unlike the main proof described above, this alternate proof does not require any computations with the Parisi formula. Instead the full RSB condition is used to ensure that low-temperature Gibbs samples have positive probability to form certain ultrametric constellations with overlap close to $1$. Existence of such constellations then forces the Hessian to have many near-zero eigenvalues. This provides a more conceptual explanation for the link between marginal stability and full RSB.

### v3

@@ 1 Introduction
${\mathcal{K}}=\left\{(\zeta,L)\in{\mathcal{N}}\times(0,\infty)\leavevmode\nobreak\ :\leavevmode\nobreak\ L>\int_{0}^{1}\zeta(s)\leavevmode\nobreak\ {\rm d}s\right\}.$ (1.4)
@@
${\mathcal{Q}}(\zeta,L)=\frac{1}{2}\left(\xi^{\prime}(0)L+\int_{0}^{1}\xi^{\prime\prime}(q)\hat{\zeta}(q)\leavevmode\nobreak\ {\rm d}q+\int_{0}^{1}\frac{{\rm d}q}{\hat{\zeta}(q)}\right)=\frac{1}{2}\left(\xi^{\prime}(1)L{-}\int_{0}^{1}\xi^{\prime\prime}(q)\Big{(}\int_{0}^{q}\zeta(s){\rm d}s\Big{)}\leavevmode\nobreak\ {\rm d}q+\int_{0}^{1}\frac{{\rm d}q}{\hat{\zeta}(q)}\right).$
@@ 1.1 Main Results
For any $\xi$ and $\varepsilon>0$, there exists $\delta>0$ such that the following holds with probability $1-e^{-cN}$ for large $N$. For all $\delta$-approximate ground states ${\boldsymbol{\sigma}}\in A_{\delta}$, the Euclidean gradient of $H_{N}$ (in ${\mathbb{R}}^{N}$) satisfies
@@
This result is stated as a proposition because a more abstract formula for $\nabla H_{N}({\boldsymbol{\sigma}})$, expressed as a derivative of the function $\xi\mapsto{\mathrm{GS}}(\xi)$, is essentially in [CS17, Remark 2] (or see [Sub24, Corollary 7]). Although we provide a detailed proof for completeness, the only novelty is an integration by parts using the stationarity conditions for $(\zeta,L)$, which yields the more tractable (1.6). Regarding the bulk spectrum (1.10), known concentration estimates imply that uniformly for all ${\boldsymbol{\sigma}}\in{\mathcal{S}}_{N}$, the spectral measure of $\nabla_{\mathrm{sp}}^{2}H_{N}({\boldsymbol{\sigma}})$ is approximately a semi-circle density with radius $2\sqrt{\xi^{\prime\prime}(1)}$ and shifted by the rescaled radial derivative $\partial_{{\mathrm{rad}}}H_{N}({\boldsymbol{\sigma}})=R({\boldsymbol{\sigma}},\nabla H_{N}({\boldsymbol{\sigma}}))$; see the proof of [Sub21, Lemma 1.3] (which is rephrased below as Lemma 1.11). Thus (1.10) follows routinely from (1.9).
@@
We say $\xi$ exhibits full RSB endpoint behavior when this condition holds. As explained in Corollary 1.14, this condition is implied by (and is presumably generically equivalent to) the usual definition of full RSB that $\zeta$ is strictly increasing in a neighborhood of $1$. We occasionally refer to the latter as strictly full RSB endpoint behavior.
When (1.11) holds, Proposition 1.2 states that the Hessian at any approximate ground state has many near-zero eigenvalues, i.e is marginally stable. Conversely when $\lambda_{+}<0$, it is natural to hope that $H_{N}$ is locally uniformly concave at near-extrema. However it is not at all obvious from existing results that the Hessian does not have a constant number of outlier eigenvalues. Our next result proves exactly this when $\xi$ contains only even degree terms.111 Theorem 1.3 and Corollary 1.4 extend with no changes when $\gamma_{1}>0$, but do require $\gamma_{p}=0$ for odd $p\geq 3$. The same holds for Corollaries 1.5 and 1.6. However deducing slow mixing from Corollary 1.6 requires modification if $\gamma_{1}>0$, as does Corollary 1.7. Indeed for the former, one must exclude the topologically trivial phase where low-temperature Langevin dynamics does mix rapidly [HS25a, Theorem 1.8]. As discussed further in Subsection 1.4, the idea is to connect extreme Hessian eigenvalues to the ground state energy of certain vector spin glasses, which can be controlled using precise interpolation bounds.
@@
Further, the local uniform concavity in (1.15) easily implies a qualitative description of the deep level sets $A_{\delta}$. As stated in Corollary 1.5 below, all connected components $A_{\delta}$ consist of small separated clusters; this is a more geometric formulation of non-marginal stability.
@@
A similar result is deduced in [BSZ20, Proposition 9.1] for a special class of $\xi$, for which a version of (1.15) follows from Kac–Rice asymptotics. We therefore give only a proof outline (just below), following their approach. As in [BSZ20] our outline uses Hessian control only at exact critical points, so it may be simplifiable.
Proof Outline of Corollary 1.5.
Consider spherical gradient-ascent flow $({\boldsymbol{\sigma}}_{t})_{t\geq 0}$ starting from any ${\boldsymbol{\sigma}}\in A_{\delta}$. This flow must have an accumulation point on the sphere by compactness, and monotonicity of $H_{N}({\boldsymbol{\sigma}}_{t})$ implies accumulation occurs at some ${\boldsymbol{\sigma}}_{*}\in A_{\delta}$. The same monotonicity also implies that ${\boldsymbol{\sigma}}_{*}$ is a critical point for $H_{N}$. Then ${\boldsymbol{\lambda}}_{1}(\nabla_{\mathrm{sp}}^{2}H_{N}({\boldsymbol{\sigma}}_{*}))\leq-C(\xi)<0$ by (1.15), so ${\boldsymbol{\sigma}}_{*}$ is a local maximum.
We now show the remaining geometric claims. Choose $C^{\prime}>0$ small compared to $C$ and $C^{\prime\prime}$ large compared to $C$, all depending only on $\xi$ (and in particular not on $\delta$). Standard smoothness estimates on $H_{N}$ (see Proposition 1.9) then imply that the intersection ${\mathcal{C}}({\boldsymbol{\sigma}}_{*})$ of $A_{\delta}$ with the ball $B_{C^{\prime}\sqrt{N}}({\boldsymbol{\sigma}}_{*})$ contains $B_{C^{\prime}\sqrt{\delta N}}({\boldsymbol{\sigma}}_{*})$ and is contained within $B_{C^{\prime\prime}\sqrt{\delta N}}({\boldsymbol{\sigma}}_{*})$. Further, ${\mathcal{C}}({\boldsymbol{\sigma}}_{*})$ is connected and thus is a connected component of $A_{\delta}$ (indeed ${\mathcal{C}}({\boldsymbol{\sigma}}_{*})$ is star-shaped relative to the center point ${\boldsymbol{\sigma}}_{*}$ since $\nabla_{\mathrm{sp}}^{2}H_{N}$ is nearly locally constant). This yields the diameter claim (modulo adjusting $C$). Since ${\boldsymbol{\sigma}}_{t}$ is a continuous path staying within $A_{\delta}$ it follows that ${\boldsymbol{\sigma}}\in{\mathcal{C}}({\boldsymbol{\sigma}}_{*})$. Since $\delta$ was small enough depending on $\xi$, we have $C/2\geq C^{\prime\prime}\sqrt{\delta}$. Hence the $C\sqrt{N}/2$-neighborhood of ${\mathcal{C}}({\boldsymbol{\sigma}}_{*})$ contains no points in $A_{\delta}\backslash{\mathcal{C}}({\boldsymbol{\sigma}}_{*})$, which yields the separation claim (modulo adjusting $C$). ∎
We pause here to emphasize two points. First, Corollary 1.5 relies crucially on the absence of Hessian outliers stated in (1.12), as even a single near-zero outlier eigenvalue could result in $A_{\delta}$ having large components. Second, the full RSB assumption is essentially necessary: when $\lambda_{+}=0$, Corollary 1.5 evidently does not hold at any ${\boldsymbol{\sigma}}\in A_{\delta}$. Indeed, it is predicted in e.g. [CK94, Section 9] that under full RSB behavior, $\delta=O(1/N)$ may suffice for $A_{\delta}$ to have components of macroscopic diameter. Rigorously, it follows (e.g. by generic perturbations as in [Pan13b, Chapter 3.7]) that when $\xi$ is full RSB near overlap $1$ in the stronger sense that $\zeta$ from (1.5) is strictly increasing on $[q,1]$, then for any $\delta>0$ and sufficiently large $N$, with high probability there exist ${\boldsymbol{\sigma}},{\boldsymbol{\sigma}}^{\prime}\in A_{\delta}$ with $R({\boldsymbol{\sigma}},{\boldsymbol{\sigma}}^{\prime})=q$.
Corollary 1.5 yields further consequences for low temperature Gibbs measures in the absence of marginal stability, which stem from similarities with the shattered phase studied in [AMS25]. In fact a substantial portion of Corollary 1.5 holds slightly more generally, in the absence of strictly full RSB endpoint behavior (defined below (1.11)). As explained in Remark 1.16 at the end of this section, this also suffices for the two corollaries below.
Using the separation of distinct components, and that low temperature Gibbs samples are approximate ground states, one obtains a plateau property for the autocorrelation function of stationary Langevin dynamics. Given an initialization ${\boldsymbol{x}}_{0}\in{\mathcal{S}}_{N}$, the spherical Langevin dynamics are given by
@@
where $P_{{\boldsymbol{x}}_{t}}^{\perp}=I_{N}-{\boldsymbol{x}}_{t}^{\otimes 2}/N$ is a rank $N-1$ projection and $\boldsymbol{B}_{t}$ is standard $N$-dimensional Browian motion. It is well known that these dynamics remain on ${\mathcal{S}}_{N}$ almost surely, with unique stationary distribution $\mu_{\beta}$. We omit the proof of the corollary below, which is exactly the same as [AMS25, Corollary 2.9]. A byproduct is exponentially slow mixing of Langevin dynamics under the same conditions (since $\mu_{\beta}$ is origin-symmetric when $\xi$ is even).
@@
The same $1/N$ scaling of $\varepsilon_{N}$ was identified for pure spherical spin glasses at very low temperature [Sub17b], and established later within the shattered phase by [AMS25]. Indeed as we explain in Subsection 2.4, Corollary 1.7 follows by the arguments in [AMS25, Section 5.1]. We note the decay condition on $\varepsilon_{N}$ is sharp: [AMS25, Proposition 2.8] shows that if $\varepsilon_{N}\leq o(1/N)$, then the expected total variation distance between $\mu_{\beta,0}$ and $\mu_{\beta,\varepsilon_{N}}$ tends to 00 in any spherical spin glass. Thus (1.17) does not hold, i.e. there is no chaos. We refer to [AMS25, Section 2.3] for further discussion, and comparison with the overlap-based notion of disorder chaos in e.g. [Cha09, CS17, CHL18, Eld20].
@@ 1.2 Connections to the Kac–Rice Formula
Whenever the Kac–Rice formula suffices to identify the ground state, the radial derivative (and hence bulk Hessian spectrum) at approximate ground states can be read off from the calculation, and the Hessians will have no outliers in the sense of Theorem 1.3; see for instance [XY24]. The reason is that the expected number of local maxima at the ground state energy level with incorrect radial derivative (or with outlier eigenvalues) is exponentially small. A similar phenomenon holds for topologically trivial spin glasses: those whose external fields are so strong that $H_{N}$ has only two critical points on ${\mathcal{S}}_{N}$, which are the global extrema [BČNS22, XZ25, HS25a]. Topological trivialization is equivalent to the minimizing $\zeta$ being the constant function 00. In general, $\zeta$ can be much more complicated; see [AZ19, AZ22b] for examples. However except for the special cases of $1$-RSB and topologically trivial models, the annealed predictions from the Kac–Rice formula are incorrect, and do not suffice to understand the behavior near the ground state (though see [Sub17b, BJ24] for applications to the geometry of positive temperature Gibbs measures). In these cases, physicists have used the replica method to predict the correct quenched critical point counts; see the recent work [FL18, KK23] as well as [BM80, CS95]. These arguments seem difficult to make rigorous at present. Despite these challenges, it is still natural to believe that exponentially rare behavior does not occur at approximate ground states, because the sets $A_{\delta}$ must be rather small by general principles of Gaussian process theory.222 For instance, [CDD13, Theorem 1.1] with $\alpha=1-\frac{\delta}{{\mathrm{GS}}(\xi)}$ shows that an independent copy $H_{N}^{\prime}$ of $H_{N}$ has typical ground state energy proportional to $\sqrt{\delta}$ when restricted to $A_{\delta}$. Via Sudakov minorization, this implies $e^{o_{\delta}(N)}$ upper bounds on the typical $\varepsilon\sqrt{N}$-covering number of $A_{\delta}$, if $\delta$ is small compared to $\varepsilon$. This intuitively suggests that $A_{\delta}$ is too small for other exponentially rare events to occur inside it (but does not constitute a proof since $A_{\delta}$ already depends on $H_{N}$). Theorem 1.3 confirms this intuition for the absence of Hessian outliers when $\xi$ is even, but uses a completely different proof technique.
@@
where $\alpha=\sqrt{\xi^{\prime\prime}+\xi^{\prime}-(\xi^{\prime})^{2}}$ and $\xi,\xi^{\prime},\xi^{\prime\prime}$ are evaluated at $1$. Of course, the exact ground state is a critical point of $H_{N}$. When $\xi$ has full RSB endpoint behavior, Proposition 1.2 implies it is marginally stable. Thus one obtains the following inequality, where the full RSB condition is necessary because ${\mathrm{GS}}(\xi)>E_{\infty}$ for pure models. By contrast ${\mathrm{GS}}(\xi)\geq E_{\infty}^{-}(\xi)$ holds for all $\xi$; see [AB13], or [HS25a, Section 7] for extensions.
@@ 1.3 On Marginal Stability
See also [MW15, FU22]. It is additionally believed that “reasonable” optimization algorithms in high-dimension rapidly reach and then get stuck in the “manifold” of marginal states [CK94, Ken24]; indeed [BAKZ22] recently conjectured that finding stable (i.e. non-marginal) local optima is computationally intractable in many disordered systems. As rigorous evidence for this, [HS24] analyzed the optimal message-passing algorithms to optimize $H_{N}$ over ${\mathcal{S}}_{N}$ and showed the resulting outputs are marginally stable outside of the topologically trivial phase, even in the more general multi-species setting. Additionally, optimal stable algorithms to optimize mean-field spin glass Hamiltonians are now understood to be closely related to full RSB [Sub21, Mon21, AMS21, HS25b, AMS23, HS23b, MZ24, JSS25]. Corollary 1.4 gives another rigorous link between full RSB and marginal stability, which however concerns low temperature statics rather than efficient algorithms.
@@ 1.4 Proof Ideas
The proofs of (1.12) and (1.13), which ensure the absence of Hessian outliers, are more delicate and given in Subsections 2.2 and 2.3 respectively. Here we perturb $H_{N}$ by considering augmented Hamiltonians of two or three replicas with constrained overlaps very close to $1$. The near-extrema for the augmented systems that one should keep in mind are obtained by starting from ${\boldsymbol{\sigma}}\in A_{\delta}$, and perturbing ${\boldsymbol{\sigma}}$ in opposite directions along a Hessian eigenvector. Thus in each case, precise estimates on the augmented ground state energy translate to bounds on the extreme eigenvalues at any approximate ground state. We find explicit interpolation parameters in the corresponding vector Parisi formulas, and perform a first-order Taylor expansion of the resulting bounds for the augmented ground state energy. This yields eigenvalue estimates which turn out to match the edge of the bulk spectrum from (1.10).
Finally Section 3 gives an alternate proof that typical Gibbs samples are marginally stable if $\beta\to\infty$ slowly with $N$, if $\xi$ is generic in addition to exhibiting full RSB endpoint behavior. This is conceptually similar to (but strictly weaker than) Corollary 1.4. Unlike the main proof described above, this alternate proof does not require any computations with the Parisi formula. Instead the full RSB condition is used to ensure that low-temperature Gibbs samples have positive probability to form certain ultrametric constellations with overlap close to $1$. Existence of such constellations then forces the Hessian to have many near-zero eigenvalues. This provides a more conceptual explanation for the link between marginal stability and full RSB.
We finally mention that the recent result [DS24, Theorem 1.4] (which appeared slightly after the present work) gives an essentially complete description of the landscape around a typical Gibbs sample in spherical spin glasses, assuming genericity and an additional finite-RSB condition. Their approach can be used to recover weaker analogs of our main results for such $\xi$ (also specialized to low temperature Gibbs samples and without a quantitative form of “high probability”). It is also worth mentioning that although the finite-RSB condition in [DS24] is fairly broad, it is in direct contrast with full-RSB and hence marginal stability.